Make New Entry,
Make Followup Entry
User name A. Tobias
Log entry time 08:09:21 on February12,2009
Entry number 260624
keyword=Owl Shift Summary
OWL SHIFT SUMMARY
00:10 Called MCC and requested nominal beam positions on the BPMs and for HARP scans (or OTR) to determine beam width (200-300microns is acceptable):
BPMA X=-1.3 Y=+2.5mm
BPMB X=-0.8 X=+2.5mm
00:55 MCC down again due to RF problem they get every hour.
01:08 Beam is back. Continue position & width adjustments.
01:35 Called MCC to verify their HARP scan & width tuning. Beam OFF due to need to reset vacuum trip points (may have to do with the hourly vacuum faults on IL04). Beam was off about 12-13min.
02:53 HARP scans look reasonable. Widths:
BPMA sigma_X=0.237mm sigma_Y=0.237mm
BPMB sigma_X=0.237mm sigma_Y=0.185mm
03:04 to 03:45 Checkout on Carbon and Pol. 3He with various rasters.
03:45 to 04:04 Set prescales for LHRS and BigBite_Buf at 14uA beam.
04:04 Start production running.
04:28 Compton and FB_A:charge_asym disconnect. Later we find out that iocha22 went down. We lost BPM and LeCroy HV readbacks. Called MCC to ask them to reset Gateway server to solve this problem.
05:44 MCC software expert had to come in to reset gateway server: iocha22: BPM and LeCroy HV readbacks are back.
07:16 MCC calls to say they have an RF problem. Took away beam.
07:22 MCC cals back and says they will send beam.
07:25 Request increase in current from 14uA to 15uA.
Checkout Run Summary w/ LHRS (unbuffered):
03:04 run 20139: checkout on Carbon/BeO, spot++ w/ 2x2.6MCC & 2uA
03:17 run 20140: checkout on Pol. 3He, spot++ w/ 2x2.6MCC & 2uA
03:26 run 20141: checkout on Pol. 3He, spot++ w/ 4x5.2MCC & 2uA
03:32 run 20142: checkout on Pol. 3He, spot++ w. 6x6.58MCC & 2uA
Production Runs w/ LHRS (unbuffered):
03:51 run 20143: production on Pol. 3He, forgot to set prescales, end run early.
04:02 run 20144: production on Pol. 3He, prescales set for 14uA beam
05:40 run 20145: production on Pol. 3He, prescales for 14uA
07:37 run 20146: production on Pol. 3He, no change in prescales
Production Runs w/ BigBite (buffered):
03:52 run 1309: production on Pol. 3He, forgot to set prescales, end run early.
04:04 run 1310: production on Pol. 3He, set prescales for 14uA beam, kill run early since ROC5 died
04:58 run 1311: production on Pol. 3He, prescales 14uA beam
05:41 run 1312: production on Pol. 3He, prescales for 14uA
06:23 run 1313: production on Pol. 3He, prescales for 14uA
07:37 run 1314: production on Pol. 3He, prescales 1 & 6 were changed